# A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems

### v2

@@ 6 Numerical experiments
In this section, we conduct a few numerical experiments to showcase the superiority and efficiency of the proposed PESS and LPESS preconditioners over the existing preconditioners to enhance the convergence speed of the Krylov subspace iterative process to solve three-by-three block SPPs. Our study involves a comparative analysis among GMRES process and the PGMRES process employing the proposed preconditioners ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS},$ and the existing preconditioners ${\mathscr{P}}_{BD}$, ${\mathscr{P}}_{IBD}$, ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS}$ and ${\mathscr{P}}_{EGSS}$ (baseline preconditioners). The numerical results are reported from the aspect of iteration counts (abbreviated as “IT”) and elapsed CPU times in seconds (abbreviated as “CPU”). Each subsystem involving $\widehat{A}$ and $\widehat{X}$ featured in Algorithm 1 is precisely solved by applying the Cholesky factorization of the coefficient matrices. For all iterative process, the initial guess vector is ${\bf u}_{0}=0\in{\mathbb{R}}^{n+m+p}$, and the termination criterion is
@@
Before proceeding, we discuss the optimal choice for the parameters of PESS preconditioner. We can rewrite the PESS preconditioner as ${\mathscr{P}}_{PESS}=s\widetilde{{\mathscr{P}}}_{PESS},$ where $\widetilde{{\mathscr{P}}}_{PESS}=\left[\begin{matrix}\frac{1}{s}\Lambda_{1}+A&B^{T}&0\\ -B&\frac{1}{s}\Lambda_{2}&-C^{T}\\ 0&C&\frac{1}{s}\Lambda_{3}\end{matrix}\right].$ Since, the pre-factor $s$ has not much effect on the performance of PESS preconditioner, investigating the optimal parameters of ${\mathscr{P}}_{PESS}$ and $\widetilde{{\mathscr{P}}}_{PESS}$ are equivalent. As we know, a preconditioner performs efficiently when it is as close as possible to the coefficient matrix $\mathcal{A},$ observe that $\widetilde{{\mathscr{P}}}_{PESS}-\mathcal{A}={\displaystyle{\frac{1}{s}\Sigma}},$ which tends to a zero matrix as $s$ tends to positive infinity for fixed $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}.$ Thus, PESS encounters enhance efficiency for large value of $s.$ However, $s$ can not be too large as the coefficient matrix $\Lambda_{1}+sA+s^{2}B^{T}(\Lambda_{2}+s^{2}C^{T}\Lambda_{3}^{-1}C)^{-1}B$ of the linear subsystem in step $3$ of Algorithm 1 becomes very-ill conditioned. Similar investigations also hold for LPESS preconditioner. Nevertheless, in Figure 7, we show the adaptability of the PESS and LPESS preconditioners by varying the parameter $s.$
@@
We consider the three-by-three block SPP (1.1) taken from [18], where
$A=\left[\begin{matrix}I\otimes G+G\otimes I&0\\ 0&I\otimes G+G\otimes I\end{matrix}\right]\in{\mathbb{R}}^{2l^{2}\times 2l^{2}},\,\,B=\left[\begin{matrix}I\otimes F&F\otimes I\end{matrix}\right]\in{\mathbb{R}}^{l^{2}\times 2l^{2}},$ and $C=E\otimes F\in{\mathbb{R}}^{l^{2}\times l^{2}},$ where $G=\frac{1}{(l+1)^{2}}\,\mathrm{tridiag}(-1,2,-1)\in{\mathbb{R}}^{l\times l},\quad F=\frac{1}{l+1}\,\mathrm{tridiag}(0,1,-1)\in{\mathbb{R}}^{l\times l}$ , and $E=\mathrm{diag}(1,l+1,\ldots,l^{2}-l+1)\in{\mathbb{R}}^{l\times l}.$ Here, $E\otimes F$ denotes the Kronecker product of two matrices $E$ and $F,$ and $\mathrm{tridiag}(a,b,c)\in{\mathbb{R}}^{l\times l}$ denotes the $l\times l$ tridiagonal matrix with diagonal, subdiagonal, and superdiagonal entries are $a,b,$ and $c,$ respectively. For this problem, the size of the matrix $\mathcal{A}$ is $4l^{2}.$
Table 2: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS and PESS PGMRES process for Example 1.
Process $l$ $16$ $32$ $48$ $64$ $80$
size($\mathcal{A}$) $1024$ $4096$ $9216$ $16384$ $25600$
@@
SS IT $4$ $4$ $4$ $4$ $4$
CPU $0.2415$ $1.3375$ $6.7059$ $33.6902$ $111.3118$
RES $087.7528e-$ 08 $085.5120E-$ 08 $084.5134E-$ 08 $083.9157E-$ 08 $083.5073E-$ 08
RSS IT $4$ $4$ $4$ $4$ $4$
CPU $0.2776$ $1.0467$ $6.5207$ $32.4428$ $110.4587$
RES $088.0898e-$ 08 $086.0033e-$ 08 $084.9839e-$ 08 $084.3510e-$ 08 $083.9097e-$ 08
EGSS IT $4$ $4$ $4$ $4$ $4$
CPU $0.3061$ $1.2771$ $6.9552$ $32.8960$ $119.7763$
RES $105.9583e-$ 10 $104.4128e-$ 10 $103.6626e-$ 10 $103.2009e-$ 10 $102.8807e-$ 10
PESS IT 2 2 2 2 2
CPU $0.2285$ $\bf 0.9854$ $\bf 4.8795$ $\bf 23.3687$ $\bf 79.9355$
$s=12$ RES $073.1630e-$ 07 $072.3239e-$ 07 $071.9486e-$ 07 $071.7260e-$ 07 $071.5754e-$ 07
LPESS IT 2 2 2 2 2
CPU $0.2834$ $\bf 0.7989$ $\bf 4.5599$ $\bf 21.1664$ $\bf 72.8376$
$s=12$ RES $073.1180e-$ 07 $072.2463e-$ 07 $071.8481e-$ 07 $071.6071e-$ 07 $071.4411e-$ 07
@@
SS IT $7$ $7$ $7$ $7$ $7$
CPU 0.2780 $1.6189$ $10.6707$ $54.4493$ $184.6025$
RES $076.7967e-$ 07 $074.9738e-$ 07 $074.1383e-$ 07 $073.6189e-$ 07 $073.2556e-$ 07
RSS IT $7$ $7$ $7$ $7$ $7$
CPU $0.3248$ $1.4644$ $10.4985$ $52.4559$ $177.1895$
RES $075.2397e-$ 07 $073.6832e-$ 07 $072.7532e-$ 07 $072.1869e-$ 07 $071.8240e-$ 07
EGSS IT $5$ $5$ $4$ $4$ $4$
CPU $0.2588$ $1.3792$ $6.7546$ $34.8358$ $118.2298$
RES $086.8491e-$ 08 $071.5239e-$ 07 $079.7099e-$ 07 $076.3466e-$ 07 $074.8159e-$ 07
PESS IT $3$ $3$ $3$ $3$ $3$
CPU $0.2971$ $1.1597$ $6.6682$ $31.2670$ $103.8616$
$s=12$ RES $087.4100e-$ 08 $087.4642e-$ 08 $087.3327e-$ 08 $086.9803e-$ 08 $086.1920e-$ 08
LPESS IT $3$ $3$ $3$ $3$ $3$
CPU $0.2226$ $0.9905$ $6.1339$ $30.0975$ $98.6443$
$s=12$ RES $091.1683e-$ 09 $092.3215e-$ 09 $093.1540e-$ 09 $093.3907e-$ 09 $093.0271e-$ 09
@@
We consider the parameter choices for the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},{\mathscr{P}}_{EGSS},{\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}$ in two cases.
  • In Case I: $\alpha=0.1$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.1,\beta=1,\gamma=0.001$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS};$ and $\Lambda_{1}=I,\Lambda_{2}=I,\Lambda_{3}=0.001I$ for ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}.$
[…]
The parameter selection in Case II is made as in [31]. The optimal value (experimentally) for $s$ in the range $[10,20]$ is determined to be $12$ for minimal CPU times. Following [18], selection of $\widehat{A}$ and $\widehat{S}$ (SPD approximations of $A$ and $S,$ respectively) for IBD preconditioner are done as follows:
@@
The test problems generated for the values of $l=16,32,48,64,80,$ and their numerical results are reported in Table 2.
@@
Figure 1: Convergence curves for IT versus RES of PGMRES process employing BD, IBD, SS, RSS, EGSS, PESS and LPESS $(s=12)$ preconditioners in Case II for Example 1.
@@
(d) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
(e) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
(f) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
(g) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=13$
(h) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=13$
Figure 2: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
@@
Figure 1 illustrates convergence curves pertaining to preconditioners BD, IBD, SS, RSS, EGSS, PESS and LPESS ($s=12$) for Case II with $l=16,32,48,64$ and $80$ These curves depict the relationship between the relative residue at each iteration step (RES) and IT counts.
To further illustrate the superiority of the PESS preconditioner, spectral distributions of $\mathcal{A}$ and the preconditioned matrices ${\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{RSS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for the Case II with $l=16$ and $s=13$ are displayed in Figure 2.
To investigate the robustness of the proposed PESS preconditioner, we measure the condition number of the ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ which is for any nonsingular matrix $A$ defined by $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2}.$ In Figure 3, the influence of the parameter $s$ on the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ is depicted. The parameter $s$ is chosen from the interval $[5,50]$ with step size one. The data presented in Table 2 and Figures 1-3 allow us to make the following noteworthy observations:
  • From Table 2, it is observed that the GMRES has a very slow convergence speed. In both Cases I and II, our proposed PESS and LPESS preconditioners outperform all other compared existing preconditioners from the aspects of IT and CPU times. For example, in Case I with $l=80,$ our proposed PESS preconditioner is $76\%$, $43\%$, $28\%$, $27\%$, and $33\%$ more efficient than the existing BD, IBD, SS, RSS, and EGSS preconditioner, respectively. Moreover, LPESS preconditioners performs approximately $78\%,$ $48\%,$ $35\%,$ $34\%,$ and $39\%$ faster than BD, IBD, SS, RSS, and EGSS preconditioner, respectively. Similar patterns are noticed for $l=16,32,48$ and $64$.
  • In Figure 1, it is evident that the PESS and LPESS preconditioners have a faster convergence speed than the other baseline preconditioners when applied to the PGMRES process for all $l=16,32,48,64$ and $80$
@@
  • According to Figure 2, the spectrum of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LESS}^{-1}\mathcal{A}$ have better clustering properties than the baseline preconditioned matrices, consequently improves the computational efficiency of our proposed PESS and LPESS PGMRES processes. Moreover, the real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ are contained in the interval $(0,0.0714]$ and for ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ are in $(0,1.9991]$ and $(0,1],$ respectively, which are consistent with bounds of Theorem 4.2, Corollaries 3.1, and 3.2, respectively. For non-real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ obtained bounds are consistent with Theorem 4.3.

### v3

@@ 7 Numerical experiments
In this section, we conduct a few numerical experiments to showcase the superiority and efficiency of the proposed PESS and LPESS preconditioners over the existing preconditioners to enhance the convergence speed of the Krylov subspace iterative process to solve three-by-three block SPPs. Our study involves a comparative analysis among GMRES process and the PGMRES process employing the proposed preconditioners ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}$ and the existing baseline preconditioners ${\mathscr{P}}_{BD}$, ${\mathscr{P}}_{IBD}$, ${\mathscr{P}}_{MAPSS}$, ${\mathscr{P}}_{SL},$ ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},$ ${\mathscr{P}}_{EGSS}$ and ${\mathscr{P}}_{RPGSS}$. The numerical results are reported from the aspect of iteration counts (abbreviated as “IT”) and elapsed CPU times in seconds (abbreviated as “CPU”). Each subsystem involving $\widehat{X}$ and $\widetilde{A}$ featured in Algorithm 1 are precisely solved by applying the Cholesky factorization of the coefficient matrices. For all iterative process, the initial guess vector is ${\bf u}_{0}=0\in{\mathbb{R}}^{n+m+p}$ and the termination criterion is
@@
Table 2: Numerical results of GMRES, BD, IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS PGMRES processes for Example 1.
Process $l$ $16$ $32$ $48$ $64$ $80$ 128
size($\mathcal{A}$) $1024$ $4096$ $9216$ $16384$ $25600$ $65536$
@@
SS IT $4$ $4$ $4$ $4$ $4$ $4$
CPU $0.2415$ $1.3375$ $6.7059$ $33.6902$ $111.3118$ $439.1221$
RES $087.7528e-$ 08 $085.5120e-$ 08 $084.5134e-$ 08 $083.9157e-$ 08 $083.5073e-$ 08 $082.7836e-$ 08
RSS IT $4$ $4$ $4$ $4$ $4$ $4$
CPU $0.2776$ $1.0467$ $6.5207$ $32.4428$ $110.4587$ $432.6755$
RES $088.0898e-$ 08 $086.0033e-$ 08 $084.9839e-$ 08 $084.3510e-$ 08 $083.9097e-$ 08 $083.1111e-$ 08
EGSS IT $4$ $4$ $4$ $4$ $4$ $4$
CPU $0.3061$ $1.2771$ $6.9552$ $32.8960$ $119.7763$ $435.0198$
RES $105.9583e-$ 10 $104.4128e-$ 10 $103.6626e-$ 10 $103.2009e-$ 10 $102.8807e-$ 10 $081.9145e-$ 08
RPGSS IT $4$ $4$ $4$ $4$ $4$ $3$
CPU $0.2249$ $1.1312$ $7.0322$ $35.0140$ $130.2593$ $371.4522$
RES $105.9497e-$ 10 $104.4042e-$ 10 $103.6494e-$ 10 $103.1838e-$ 10 $102.8603e-$ 10 $079.9326e-$ 07
PESS IT 2 2 2 2 2 $\bf{2}$
CPU $0.2285$ $\bf 0.9854$ $\bf 4.8795$ $\bf 23.3687$ $\bf 79.9355$ $\bf 283.8561$
$s=12$ RES $073.1630e-$ 07 $072.3239e-$ 07 $071.9486e-$ 07 $071.7260e-$ 07 $071.5754e-$ 07 $071.3135e-$ 07
LPESS IT 2 2 2 2 2 $\bf 2$
CPU $0.2834$ $\bf 0.7989$ $\bf 4.5599$ $\bf 21.1664$ $\bf 72.8376$ $\bf 268.3114$
$s=12$ RES $073.1180e-$ 07 $072.2463e-$ 07 $071.8481e-$ 07 $071.6071e-$ 07 $071.4411e-$ 07 $071.1441e-$ 07
@@
SS IT $7$ $7$ $7$ $7$ $7$ $7$
CPU 0.2780 $1.6189$ $10.6707$ $54.4493$ $184.6025$ $703.1082$
RES $076.7967e-$ 07 $074.9738e-$ 07 $074.1383e-$ 07 $073.6189e-$ 07 $073.2556e-$ 07 $072.5956e-$ 07
RSS IT $7$ $7$ $7$ $7$ $7$ $7$
CPU $0.3248$ $1.4644$ $10.4985$ $52.4559$ $177.1895$ $690.1611$
RES $075.2397e-$ 07 $073.6832e-$ 07 $072.7532e-$ 07 $072.1869e-$ 07 $071.8240e-$ 07 $071.2720e-$ 07
EGSS IT $5$ $5$ $4$ $4$ $4$ $4$
CPU $0.2588$ $1.3792$ $6.7546$ $34.8358$ $118.2298$ $439.7907$
RES $086.8491e-$ 08 $071.5239e-$ 07 $079.7099e-$ 07 $076.3466e-$ 07 $074.8159e-$ 07 $074.1376e-$ 07
RPGSS IT $4$ $4$ $4$ $4$ $4$ $3$
CPU $0.2445$ $1.2987$ $8.1982$ $28.5129$ $104.9127$ $397.4766$
RES $085.2642e-$ 08 $071.1366e-$ 07 $087.1116e-$ 08 $078.2780e-$ 07 $075.6396e-$ 07 $072.5131e-$ 07
PESS IT $3$ $3$ $3$ $3$ $3$ $3$
CPU $0.2971$ $1.1597$ $6.6682$ $31.2670$ $103.8616$ $371.1214$
$s=12$ RES $087.4100e-$ 08 $087.4642e-$ 08 $087.3327e-$ 08 $086.9803e-$ 08 $086.1920e-$ 08 $084.0967e-$ 08
LPESS IT $3$ $3$ $3$ $3$ $3$ $3$
CPU $0.2226$ $0.9905$ $6.1339$ $30.0975$ $98.6443$ $359.8141$
$s=12$ RES $091.1683e-$ 09 $092.3215e-$ 09 $093.1540e-$ 09 $093.3907e-$ 09 $093.0271e-$ 09 $091.6141e-$ 09
@@
Problem formulation: We consider the three-by-three block SPP (1.1) taken from [8] with
$A=\left[\begin{matrix}I\otimes G+G\otimes I&0\\ 0&I\otimes G+G\otimes I\end{matrix}\right]\in{\mathbb{R}}^{2l^{2}\times 2l^{2}},\,\,B=\left[\begin{matrix}I\otimes F&F\otimes I\end{matrix}\right]\in{\mathbb{R}}^{l^{2}\times 2l^{2}},$ $C=E\otimes F\in{\mathbb{R}}^{l^{2}\times l^{2}},$ where $G=\frac{1}{(l+1)^{2}}\,\mathrm{tridiag}(-1,2,-1)\in{\mathbb{R}}^{l\times l},\quad F=\frac{1}{l+1}\,\mathrm{tridiag}(0,1,-1)\in{\mathbb{R}}^{l\times l}$ and $E=\mathrm{diag}(1,l+1,\ldots,l^{2}-l+1)\in{\mathbb{R}}^{l\times l}.$ Here, $E\otimes F$ denotes the Kronecker product of two matrices $E$ and $F$ and $\mathrm{tridiag}(a,b,c)\in{\mathbb{R}}^{l\times l}$ denotes the $l\times l$ tridiagonal matrix with diagonal entries $b$, subdiagonal entries $a$, and superdiagonal entries $c$. For this problem, the size of the matrix $\mathcal{A}$ is $4l^{2}.$
Parameter selection: Following [8], selection of $\widehat{A}$ and $\widehat{S}$ (SPD approximations of $A$ and $S,$ respectively) for IBD preconditioner are done as follows:
@@
For the preconditioner ${\mathscr{P}}_{\textit{MAPSS}},$ we take $\alpha=\sqrt[4]{\frac{tr(BB^{T}C^{T}C)}{m}}$ and $\beta=10^{-4}$ as in [21]. We consider the parameter choices for the preconditioners ${\mathscr{P}}_{SS},{\mathscr{P}}_{RSS},{\mathscr{P}}_{EGSS},$ ${{\mathscr{P}}_{RPGSS}},$ ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}$ in the following two cases.
  • In Case I: $\alpha=0.1$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=0.1,\beta=1,\gamma=0.001$ and $P=I,Q=I,W=I$ for ${\mathscr{P}}_{EGSS}$ and ${\mathscr{P}}_{RPGSS}$; and $\Lambda_{1}=I,\Lambda_{2}=I,\Lambda_{3}=0.001I$ for ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}.$
  • In Case II: $\alpha=1$ for ${\mathscr{P}}_{SS}$ and ${\mathscr{P}}_{RSS};$ $\alpha=1,\beta=1,\gamma=0.001$ and $P=A,Q=I,W=CC^{T}$ for ${\mathscr{P}}_{EGSS}$ and ${\mathscr{P}}_{RPGSS}$; and $\Lambda_{1}=A,\Lambda_{2}=I,\Lambda_{3}=0.001CC^{T}$ for ${\mathscr{P}}_{PESS}$ and ${\mathscr{P}}_{LPESS}.$
The parameter selection in Case II is made as in [39].
Results for experimentally found optimal parameter: The optimal value (experimentally) for $s$ in the range $[10,20]$ is determined to be $12$ for minimal CPU times. The test problems generated for the values of $l=16,32,48,64,80,{128},$ and their numerical results are reported in Table 2.
Table 3: Numerical results of PESS-I, LPESS-I, PESS-II and LPESS-II PGMRES processes for Example 1.
Process $l$ $16$ $32$ $48$ $64$ $80$ $128$
size($\mathcal{A}$) $1024$ $4096$ $9216$ $16384$ $25600$ $65536$
PESS-I IT $2$ $2$ $2$ $2$ $2$ $2$
($s=1,$ $\Lambda_{1}=0.01I,$ CPU $0.2517$ $0.8373$ $4.6626$ $21.8133$ $77.9420$ $308.1418$
$\Lambda_{2}=0.1I,$ $\Lambda_{3}=0.001I$) RES $074.4970e-$ 07 $073.8763e-$ 07 $073.6261e-$ 07 $073.4880e-$ 07 $073.3983e-$ 07 $077.7830e-$ 07
LPESS-I IT $2$ $2$ $2$ $2$ $2$ $2$
($s=1,$ $\Lambda_{2}=0.1I,$ CPU $0.3070$ $0.8184$ $4.7244$ $21.6534$ $82.3687$ $288.9701$
$\Lambda_{3}=0.001I$ ) RES $073.1116e-$ 07 $072.2410e-$ 07 $071.8435e-$ 07 $071.6030e-$ 07 $071.4375e-$ 07 $071.1441e-$ 07
PESS-II IT $3$ $3$ $3$ $3$ $3$ $3$
($s=s_{est},$ $\Lambda_{1}=A,$ CPU $0.20373$ $1.05327$ $5.84150$ $29.3914$ $104.9663$ $424.2004336$
$\Lambda_{2}=\beta_{est}I,$ $\Lambda_{3}=10^{-4}CC^{T}$) RES $082.2013e-$ 08 $083.4850e-$ 08 $084.3689e-$ 08 $085.0705e-$ 08 $085.8005e-$ 08 $087.2906e-$ 08
LPESS-II IT $3$ $3$ $3$ $3$ $3$ $3$
($s=s_{est},$ $\Lambda_{2}=\beta_{est}I,$ CPU $0.19215$ $1.04964$ $4.63143$ $24.20924$ $102.84651$ $389.10761$
$\Lambda_{3}=10^{-4}CC^{T}$) RES $091.8126e-$ 09 $131.0207e-$ 13 $158.0620e-$ 15 $149.1601e-$ 14 $135.7292e-$ 13 $071.1950e-$ 07
Results using parameters selection strategy in Section 6: To demonstrate the effectiveness of the proposed preconditioners PESS and LPESS using the parameters discussed in Section 6, we present the numerical results by choosing $s=1,$ $\Lambda_{1}=0.01I,\Lambda_{2}=0.1I,\Lambda_{3}=0.001I$ (denoted by PESS-I and LPESS-I) and $s=s_{est},$ $\Lambda_{1}=A,$ $\Lambda_{2}=\beta_{est}I$ and $\Lambda_{3}=10^{-4}C^{T}C$ (denoted by PESS-II and LPESS-II) for the proposed preconditioner. These results are summarized in Table 3.
Convergence curves: Figure 1 illustrates convergence curves pertaining to preconditioners BD, IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS ($s=12$) for Case II with $l=16,32,48,64,80$ and $128$. These curves depict the relationship between the relative residue at each iteration step (RES) and IT counts.
Spectral distributions: To further illustrate the superiority of the PESS preconditioner, spectral distributions of $\mathcal{A}$ and the preconditioned matrices ${\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},$ ${{\mathscr{P}}_{MAPSS}^{-1}\mathcal{A}},$ ${{\mathscr{P}}_{SL}^{-1}\mathcal{A}},$ ${\mathscr{P}}_{SS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A},$ ${{\mathscr{P}}_{RPGSS}^{-1}\mathcal{A}},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for the Case II with $l=16$ and $s=13$ are displayed in Figure 2.
@@
(f) $l=128$
Figure 1: Convergence curves for IT versus RES of PGMRES processes employing BD, IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS $(s=12)$ preconditioners in Case II for Example 1.
@@
(d) ${\mathscr{P}}_{MAPSS}^{-1}\mathcal{A}$
(e) ${\mathscr{P}}_{SL}^{-1}\mathcal{A}$
(f) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
(g) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
(h) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
(i) ${\mathscr{P}}_{RPGSS}^{-1}\mathcal{A}$
(j) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=13$
(k) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=13$
Figure 2: Spectral distributions of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{{\mathscr{P}}_{MAPSS}^{-1}\mathcal{A},{\mathscr{P}}_{SL}^{-1}\mathcal{A},}{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{{\mathscr{P}}_{RPGSS}^{-1}\mathcal{A},}{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
(a) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=13$
(b) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=13$
Figure 3: Spectral bounds for ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with $l=16$ for Example 1.
Spectral bounds: Furthermore, we draw the eigenvalue bounds of Theorems 4.1 and 4.2 for the PESS preconditioned matrix. The $|\lambda-1|=1$ of Theorem 4.1 is drawn by the green unit circle in Figure 3(a) and the points in blue color are the bounds of Theorem 4.2. To draw the eigenvalues bounds in Theorem 5.1 for LPESS preconditioned matrix, we define the following circles:
$C_{1}:=\left\{\lambda\in\mathbb{C}:|\lambda|=\sqrt{\frac{\tilde{\theta}_{\max}}{1+s\vartheta_{\min}+s^{2}\tilde{\theta}_{\max}}}\right\},\leavevmode\nobreak\ C_{2}:=\left\{\lambda\in\mathbb{C}:|\lambda|=\frac{\vartheta_{\min}}{2+s\vartheta_{\min}}\right\},$
$C_{3}:=\left\{\lambda\in\mathbb{C}:\left|\lambda-\frac{1}{s}\right|=\frac{\vartheta_{\min}}{2+s\vartheta_{\min}}\right\}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ C_{4}:=\left\{\lambda\in\mathbb{C}:\left|\lambda-\frac{1}{s}\right|=\frac{1}{s(1+s\sqrt{\tilde{\theta}_{\max}})}\right\}.$
In Figure 3(b), $C_{1}$ is in red color, $C_{2}$ is in blue color, $C_{3}$ is in black color and $C_{4}$ is in green color. Additionally, the eigenvalue bounds in Theorems 4.2, 4.3 and 5.1 are also presented in Table 4.
Table 4: Spectral bounds for ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for case II with $l=16$ for Example 1.
Real eigenvalue $\lambda$ Non-real eigenvalue $\lambda$
Bounds of Theorem 4.2 Bounds of Theorem 4.3
${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ contained in $0.0667\leq|\lambda|\leq 0.0739$
the interval $4.258\times 10^{-5}\leq\Re(\lambda/(1-s\lambda))\leq 0.5$
$(0,0.071429]$ $|\Im(\lambda/(1-s\lambda))|\leq 31.6386$
Bounds of Theorem 5.1 Bounds of Theorem 5.1
${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ contained in $0.0285\leq|\lambda|\leq 0.07693$
the interval $1.8666\times 10^{-4}\leq|\lambda-\frac{1}{s}|\leq 0.04384$
$(0.0416,0.0769]$
Condition number analysis: To investigate the robustness of the proposed PESS preconditioner, we measure the condition number of the ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ which is for any nonsingular matrix $A$ is defined by $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2}.$ In Figure 4, the influence of the parameter $s$ on the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ for Case II with $l=32$ is depicted. The parameter $s$ is chosen from the interval $[5,50]$ with step size one.
Discussions: The data presented in Tables 2 and 3 and Figures 1-4 allow us to make the following noteworthy observations:
  • From Table 2, it is observed that the GMRES has a very slow convergence speed. In both Cases I and II, our proposed PESS and LPESS preconditioners outperform all other compared existing preconditioners from the aspects of IT and CPU times. For example, in Case I with $l=80,$ our proposed PESS preconditioner is $76\%$, $43\%$, $36\%,$ $22\%,$ $28\%$, $27\%,$ $33\%$ and $39\%$ more efficient than the existing BD, IBD, MAPSS, SL, SS, RSS, EGSS and RPGSS preconditioner, respectively. Moreover, LPESS preconditioners performs approximately $78\%,$ $48\%,$ $38\%,$ $29\%$, $35\%,$ $34\%,$ $39\%$ and $44\%$ faster than BD, IBD, MAPSS, ${\mathscr{P}}_{SL},$ SS, RSS, EGSS and RPGSS preconditioner, respectively. Similar patterns are noticed for $l=16,32,48,64$ and $128$.
  • Comparing the results of Tables 2 and 3, we observe that in both cases, PESS and LPESS preconditioners outperform the existing baseline preconditioners and IT are the same as in the experimentally found optimal parameters. This shows that the parameter selection strategy in Section 6 is effective.
  • In Figure 1, it is evident that the PESS and LPESS preconditioners have a faster convergence speed than the other baseline preconditioners when applied to the PGMRES process for all $l=16,32,48,64,80$ and $128$
@@
  • According to Figure 2, the spectrum of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LESS}^{-1}\mathcal{A}$ have better clustering properties than the baseline preconditioned matrices, consequently improves the computational efficiency of our proposed PESS and LPESS PGMRES processes. Moreover, the real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ are contained in the interval $(0,0.071429]$ and for ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$ are in $(0,1.9991]$ and $(0,1],$ respectively, which are consistent with bounds of Theorem 4.2, Corollaries 3.1 and 3.2, respectively. For non-real eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A},$ obtained bounds are consistent with Theorem 4.3. Moreover, from Figure 3(b), we observe that eigenvalues of ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ are contained in $C_{2}\leq\lambda\leq C_{1}\cap C_{4}\leq\lambda\leq C_{4}$, which shows the consistency of the bounds in Theorem 5.1.